OpenStudy (anonymous):

how do i solve y=x^2-4

OpenStudy (anonymous):

x=-2 and x=2

OpenStudy (anonymous):

is a sum of differences equation so it simplifies to (x-2)(x+2)

OpenStudy (anonymous):

and then find the zeros x-2+0 and x+2+0

jhonyy9 (jhonyy9):

us formula a^2 -b^2 =(a-b)(a+b)

jhonyy9 (jhonyy9):

and you will get from your equation because you have there y=x^2 -2^(2) what is equal with (x-2)(x+2) - so in this case you have y=(x-2)(x+2) (x-2)(x+2)=0 ---- x-2=0 --- x=2 x+2=0 --- x=-2